Find the LCD for the following; then add as indicated.
1/4+1/3+1/5+1/6

Answer by RAY100(1637)

1/3+1/4+1/5+1/6
.
lcd = 60 (3*2*2*5)
,
{20+15+12+10) / 60
.
57 /60
.
check,,1/3 +1/4 +1/5+1/6 = .33+.25+.2+.167 =.947 = 57/60,,,ok